# Large Differences Between Stochastic and Deterministic Kinetics in a Simple Autocatalytic Reaction Network

### v2

@@ IV Summary and Discussion
In this work, we examine differences between deterministic and stochastic kinetics of the Finke-Watkzy model (FWM): irreversible autocatalysis, $\mathrm{A}+\mathrm{B}\rightarrow 2\mathrm{B}$ (2), paired with an irreversible (pseudo)first-order reaction, $\mathrm{A}\rightarrow\mathrm{B}$ (1). The FWM can be used to describe the formation of transition metal colloids in solution, protein misfolding followed by protein aggregation, and the spread of a non-fatal infectious disease that can be transmitted by direct or indirect contact.
@@
The relatively large difference between the time dependence of the average number of molecules and the time dependence of the concentration of the same chemical species has already been reported for the autocatalytic reaction (2) and termed ’stochastic delay’ Schuster (2019). By ’relatively large’ we mean large compared to a simple non-autocatalytic irreversible bimolecular reaction $\mathrm{A}+\mathrm{B}\rightarrow 2\mathrm{C}$ for a system with the same total number of molecules. Here we show that the presence of the second, parallel first-order reaction (apart from the autocatalysis) can greatly amplify the discrepancies between stochastic and deterministic kinetics. This is the case when there are initially no B molecules and the rate constant $k_{1}$ for the first-order reaction (1) is much smaller than the rate constant $k_{2}$ for the autocatalytic reaction (2). The fluctuations around the average number of A molecules are also very large in this case.
@@ Appendix B The derivation of the expressions for the first and second moments of $P_{n}(t)$
${\Gamma(z-n)}=(-1)^{n-1}\frac{\Gamma(-z)\Gamma(1+z)}{\Gamma(n+1-z)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ n\in\mathbb{Z},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ z\notin\mathbb{Z},$ (88)
@@
${\Gamma(z)\Gamma(1-z)}=\frac{\pi}{\sin(\pi z)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ z\notin\mathbb{Z}.$ (89)
@@ C.2 Reversible generalization of the Finke-Watkzy model (3)–(4)
$=$ $\leavevmode\nobreak\ \leavevmode\nobreak\ (\mathcal{K}^{+}_{2}+\mathcal{K}^{-}_{2})(a-a_{+})(a-a_{-}),$

### v3

@@ IV Summary and Discussion
In this work, we examine differences between deterministic and stochastic kinetics of the Finke-Watkzy model (FWM): irreversible autocatalysis, $\mathrm{A}+\mathrm{B}\rightarrow 2\mathrm{B}$ (2), paired with an irreversible (pseudo)first-order reaction, $\mathrm{A}\rightarrow\mathrm{B}$ (1).
@@
The relatively large difference between the time dependence of the average number of molecules and the time dependence of the concentration of the same chemical species has already been reported for the autocatalytic reaction (2) and termed ’stochastic delay’ Schuster (2019). By ’relatively large’ we mean large compared to a simple non-autocatalytic irreversible bimolecular reaction $\mathrm{A}+\mathrm{B}\rightarrow 2\mathrm{C}$ for a system with the same total number of molecules. Here we show that the presence of the second, parallel first-order reaction can greatly amplify the discrepancies between stochastic and deterministic kinetics. This is the case when there are initially no B molecules and the rate constant $k_{1}$ for the first-order reaction (1) is much smaller than the rate constant $k_{2}$ for the autocatalytic reaction (2). The fluctuations around the average number of A molecules are also very large in this case.
@@ Appendix B The derivation of the expressions for the first and second moments of $P_{n}(t)$
${\Gamma(z-n)}=(-1)^{n-1}\frac{\Gamma(-z)\Gamma(1+z)}{\Gamma(n+1-z)},\penalty 10000\ \penalty 10000\ \penalty 10000\ n\in\mathbb{Z},\penalty 10000\ \penalty 10000\ \penalty 10000\ z\notin\mathbb{Z},$ (82)
@@
${\Gamma(z)\Gamma(1-z)}=\frac{\pi}{\sin(\pi z)},\penalty 10000\ \penalty 10000\ \penalty 10000\ z\notin\mathbb{Z}.$ (83)
@@ C.2 Reversible generalization of the Finke-Watkzy model (3)–(4)
$=$ $\penalty 10000\ \penalty 10000\ (\mathcal{K}^{+}_{2}+\mathcal{K}^{-}_{2})(a-a_{+})(a-a_{-}),$